the pressure of a fluid is a linear function of of volume and the internal energy of the fluid is U=34+3PV (s.i.units).then find a,b,w,change in internal energy and q for vhange in state from(100pa,3m^3) to (400pa,6m^3)

Grade:12

1 Answers

Gopesh Kumar IIT Roorkee
9 years ago
The equation for pressure is not given. So let it be P=a+bV using the given data 100=a+ 3b and 400=a+6b solving we get, b=100 and a= -200. Now change in internal energy = 3(300-2400)=-6300J work done= -200(3)+50*27 =600- 10350 J = -9750J So, Q= U-PV = 3450J